OpenStudy (anonymous):

Ben and Chris start at the same time from places 96 miles apart to cycle toward one another, Ben at 16 mi/h and Christ at 12 mi/h. If Ben rests for 30 minutes on the way, after how many hours will they meet?

OpenStudy (anonymous):

...

OpenStudy (dumbcow):

d = r*t --> 16(t-.5) + 12t = 96 28t -8 = 96 28t = 104 t = 3.714

OpenStudy (anonymous):

i think we should not subtract 0.5 but add, since it's part of the travel. . ?

OpenStudy (anonymous):

Is it?

OpenStudy (anonymous):

 rest is part of the travel. even Ben is resting for 30 mins. , Chris is still moving. . therefor, the rest is still part of travel..